# A try for dark energy in quantum field theory: The vacuum energy of neutrino field

### v2

@@ IV Conclusion and Discussion
In this paper, the vacuum energy of a field has been investigated, which is another UV divergence of power-law type in QFT. The active vacuum energy of neutrino fields can naturally produce the required dark energy density in the ripple model, and hence a prediction of a neutrino with a typical mass around 10 meV should be present. Considering the naturalness at an energy scale $\mu_{\Lambda}$, the normal hierarchical spectrum of neutrinos is preferred, and there is a mass-energy window for the neutrino mass spectrum and the dark energy density. The neutrino mass window set by dark energy can be examined by the future experiments. In addition, the value of the energy scale $\mu_{\Lambda}$ describing the active contribution of a field here is derived by the observation of dark energy density, and it is taken as an input parameter with naturalness. If this scale is considered from a theoretical point of view, it may be related to the inflationary field, or the origin of the neutrino mass, etc. Moreover, for UV-divergence inputs in QFT, there may be another route besides the routes mentioned in this article. All roads lead to the same goal. With the joint explorations of multiple routes, we hope it will be possible to get close to the harmonious and unified physical structure hidden beneath UV divergences. We look forward to more explorations.
@@ Acknowledgements.
The author would like to thank James M. Cline for helpful discussions. This work was partly supported by the fund of SWUST under the contract no. 23cxtd59.

### v4

@@ IV Conclusion and Discussion
In this paper, the vacuum energy’s power-law divergence has been established. The physical result in the UV-free scheme indicates that it is not a problem in the UV region. Freed from the UV divergence covering, the actual question is that contributions of heavy fields should be suppressed by unknown mechanisms. Here we give more interpretations about two common questions as mentioned by Ref. Cline:2024klo , i.e., why we choose the value of $\mu_{\Lambda}$ in meV scale, and why we ignore the heavy fields that could contribute. To the first question about the value of $\mu_{\Lambda}$, it is currently lacking pathways to derive its specific value from fundamental principles (e.g., the values of particle masses and coupling constants in SM). Here the scale $\mu_{\Lambda}$ is taken as an input parameter, and its parameter space is derived by the observation of dark energy density with naturalness. This energy scale preserves the spacetime flatness of the present-day universe. Theoretically, this scale may link to the inflationary field or neutrino mass origin, etc. To the second question, this work explores the $\mu_{\Lambda}-$characterized scale decoupling of vacuum energy from microscopic quantity to macroscopic effect, implying heavy fields remain inert. The answer to it is a long exploring journey.
To account for the dark energy density, a neutrino with a typical mass around 10 meV is expected in the ripple description. Considering the naturalness at the scale $\mu_{\Lambda}$, the normal hierarchical spectrum of neutrinos is preferred, and there is a mass-energy window for the neutrino mass spectrum and the dark energy density. The neutrino mass window set by dark energy can be examined by the future experiments. Moreover, for UV-divergence inputs in QFT, there may be another route besides the routes mentioned in this article. All roads lead to the same goal. With the joint explorations of multiple routes, we hope it will be possible to get close to the harmonious and unified physical structure hidden beneath UV divergences. We look forward to more explorations.
@@ Acknowledgements.
The author thanks James M. Cline for useful discussions. This work was partly supported by the fund of SWUST under the contract no. 23cxtd59.